# Revisiting Multivariate Time Series Forecasting with Missing Values

### v2

@@ Abstract
Missing values are common in real-world time series, and multivariate time series forecasting with missing values (MTSF-M) has become a crucial area of research for ensuring reliable predictions. To address the challenge of missing data, current approaches have developed an imputation-then-prediction framework that uses imputation modules to fill in missing values, followed by forecasting on the imputed data. However, this framework overlooks a critical issue: there is no ground truth for the missing values, making the imputation process susceptible to errors that can degrade prediction accuracy. In this paper, we conduct a systematic empirical study and reveal that imputation without direct supervision can corrupt the underlying data distribution and actively degrade prediction accuracy. To address this, we propose a paradigm shift that moves away from imputation and directly predicts from the partially observed time series. We introduce Consistency-Regularized Information Bottleneck (CRIB), a novel framework built on the Information Bottleneck principle. CRIB combines a unified-variate attention mechanism with a consistency regularization scheme to learn robust representations that filter out noise introduced by missing values while preserving essential predictive signals. Comprehensive experiments on four real-world datasets demonstrate the effectiveness of CRIB, which predicts accurately even under high missing rates. Our code implementation is available in https://github.com/Muyiiiii/CRIB.
@@ 1 Introduction
Multivariate time series forecasting (MTSF), which aims to predict future values of multiple variates based on historical observations, plays an important role in many domains, such as traffic flow forecasting (Shang et al., 2022; Yu et al., 2017; Bai et al., 2020), financial analysis (Schaffer et al., 2021; Zivot & Wang, 2006; Hu et al., 2025b; a), and weather prediction (Zheng et al., 2015; Wu et al., 2021; Tan et al., 2022). However, due to uncontrollable factors such as data collection difficulties and transmission failures (Li et al., 2023; Marisca et al., 2022; Cini et al., 2021; Zhang et al., 2025a), real-world multivariate time series data is often partially observed, with missing values scattered throughout the series. These missing values inevitably introduce noise, leading to distribution shifts and disrupting the variate correlations. MTSF models (Cao et al., 2020; Liu et al., 2022; Ekambaram et al., 2023; Hu et al., 2025e), which typically rely on complete data, are highly sensitive to such shifts and correlation destruction, thus failing to make accurate predictions (Zhou et al., 2023; Hu et al., 2025c). This has driven increasing interest in multivariate time series forecasting with missing values (MTSF-M) (Cao et al., 2018; Zuo et al., 2023; Tang et al., 2020), where the objective is to generate accurate and robust forecasts despite the presence of incomplete data.
@@
To mitigate the impact of missing values, recent MTSF-M research (Yu et al., 2025; Peng et al., 2025) has focused on enhancing observed data by imputing missing values to improve prediction performance. One common approach is the two-stage framework, where an imputation module (Wu et al., 2022; Cao et al., 2018; Du et al., 2023) first fills in the missing values, and a forecasting model then predicts future values based on the imputed data (Peng et al., 2025; Chen et al., 2023; Wu et al., 2015). Moreover, to reduce error accumulation between these two stages of two separate models, some studies have proposed an end-to-end framework (Yu et al., 2024; 2025) that imputes missing values progressively during encoding and performs forecasting using the imputed representations. Overall, these methods generally follow an imputation-then-prediction paradigm, aiming to improve forecasting accuracy by mitigating the negative effects of missing values compared to directly applying forecasting models to incomplete data.
@@
Figure 1: Analysis of the imputation-then-prediction paradigm on PEMS-BAY (40% missing rate). (a) t-SNE visualizations show that current imputation modules cannot recover the original data distribution and their forecasts mismatch with the prediction target, while our direct-prediction method aligns better with the target. (b, c) Correlation maps reveal that imputation fails to recover true variate correlations, whereas our method preserves underlying correlations more effectively.
@@
However, current MTSF-M methods ignore a critical limitation in real-world applications: there is no ground truth for missing values. In such scenarios, the imputation module of the current MTSF-M methods would lack reliable guidance, which means the imputed values and reconstructed correlations cannot be guaranteed to be accurate with only the final prediction guidance. As a result, noise would propagate into the prediction stage and degrade forecasting performance, particularly when the missing rate is high. To investigate this issue, we conduct an empirical analysis of representative imputation-then-prediction methods, where original and observed data denote the complete and partially observed data, respectively. This includes the two-stage framework combining TimesNet (Wu et al., 2022) for imputation and DLinear (Zeng et al., 2023) for forecasting, as well as the end-to-end framework BiTGraph (Chen et al., 2023). Fig. 1 illustrates the empirical results, where panel (a) visualizes the distributions of imputed and predicted values, and panels (b) and (c) present the correlations among variates. Our findings highlight two key phenomena:
@@
  • Improper imputation can corrupt the observed data. Current MTSF-M frameworks commonly employ imputation modules to recover missing values. However, as shown in Fig. 1 (a-1, b), without enough direct supervision, imputed values deviate considerably from the distribution of the original complete data, and the underlying correlations among variates are not correctly reconstructed. The deterioration of both the data distribution and variate correlations suggests that imputation with only prediction guidance can degrade the observed data rather than repair it.
@@
  • Flawed imputation, in turn, leads to poor prediction performance. Errors from the imputation stage inevitably propagate into forecasting. As shown in Fig. 1 (a-2, c), the predictions exhibit large deviations from the prediction targets. Notably, even a simple model DLinear applied directly to incomplete observed data outperforms a more complex framework that combines TimesNet for imputation with DLinear for prediction. These findings indicate that a flawed imputation stage can actively harm, rather than enhance, the forecasting capabilities of a model.
@@
  • Empirical analysis: We perform a systematic empirical analysis of the dominant imputation-then-prediction paradigm for MTSF-M. We reveal that, guided only by a prediction objective, imputation modules can corrupt the observed data distribution and degrade prediction performance.
  • Method: We propose a novel direct-prediction method, CRIB, which removes the imputation completely. CRIB is an IB-based method that integrates a unified-variate attention mechanism and consistency regularization to get refined representations, effectively balancing the tradeoff between filtering out noise and preserving task-relevant signals.
  • Experiments: We conduct comprehensive experiments on four real-world benchmarks and show that CRIB significantly outperforms existing state-of-the-art methods by an average of 18%, especially under high missing rates. Our results validate the superiority of the proposed direct-prediction approach over the imputation-then-prediction paradigm.
@@ 2 Preliminaries
IB theory (Tishby & Zaslavsky, 2015; Voloshynovskiy et al., 2019) provides an information-theoretic framework for learning compact and informative representations. Given the partially observed input $X^{\text{o}}$ and the prediction target $Y$, the goal is to learn a latent representation $Z$ that is maximally compressive with respect to $X^{\text{o}}$ while remaining maximally informative about $Y$. This trade-off in CRIB is formalized by the following objective:
@@
Here, $\theta$ represents the learnable parameters of our proposed CRIB. $I(Z;X^{\text{o}})$ and $I(Y;Z)$ are the mutual information terms measuring compactness and informativeness, respectively. The Lagrange multiplier $\beta\in\mathbb{R}$ controls the balance between these two terms (Tishby et al., 2000). Furthermore, under standard assumptions in the IB literature (Alemi et al., 2016; Chalk et al., 2016; Ma et al., 2023), the joint distribution of the variables can be factorized as:
$p(X^{\text{o}},Y,Z)=p(Z|X^{\text{o}},Y)p(Y|X^{\text{o}})p(X^{\text{o}})=p(Z|X^{\text{o}})p(Y|X^{\text{o}})p(X^{\text{o}}),$ (2)
@@ 3 Methodology
Figure 2: Overall framework of CRIB. (a) Data Augmentation creates a more challenging view of the partially observed data $X^{\text{o}}$ by generating an augmented version $X^{\text{Aug}}$. (b) The Patching Embedding layer converts the $X^{\text{o}}$ and $X^{\text{Aug}}$ into robust patch-level feature representations $H$ and $H^{\text{Aug}}$. (c) The Unified-Variate Attention mechanism models the global correlations between all the patches within $H$ and $H^{\text{Aug}}$ to produce refined representations $Z$ and $Z^{\text{Aug}}$. (d) Consistency Regularization aligns the representations from the original $Z$ and the augmented views $Z^{\text{Aug}}$. The entire process is guided by the IB principles of compactness and informativeness to produce the final forecast $\widehat{Y}$.
As illustrated in Fig. 2, our proposed model, CRIB, bypasses the problematic imputation stage by performing forecasts directly on the partially observed data. The architecture is composed of several key stages, each designed to address the challenges of learning from partially observed data. First, to handle the raw, sparse input, we introduce a Patching Embedding layer that employs a Temporal Convolutional Network (TCN) (Bai et al., 2018) to learn robust local feature representations from available data points. Second, to capture the complex global correlations that are disrupted by missingness, a Unified-Variate Attention mechanism models correlations across all patches simultaneously. Third, to ensure the model learns features that are stable and invariant to different missingness, especially under high missing rates, we introduce a Consistency Regularization scheme based on data augmentation. The entire learning process is guided by the IB principle, which provides a theoretical foundation for learning a representation that is maximally compressive against noise while being sufficiently informative for the forecasting task.
@@ 3.1 Patching Embedding
$\text{TE}(t,m)=\begin{cases}\ \sin(t/10000^{2t/P})&\text{ if }m=2k,\\ \ \cos(t/10000^{2t/P})&\text{ if }m=2k+1,\end{cases}$ (3)
@@ 3.4 Information Bottleneck Guidance
To enhance the quality of the learned representations $Z$ and improve forecasting accuracy, we propose an IB-based guidance. This guidance aims to balance compactness (filtering out irrelevant information) with informativeness (preserving relevant task-specific signals), allowing CRIB to focus on the most significant factors for accurate forecasting. In this section, we present how the compactness and informativeness principles are formulated and implemented in our framework. Full derivations are detailed in Appendix A.
@@ 3.4.1 Compactness Principle
The compactness principle, which aims to minimize the mutual information $I_{\theta}(Z;X^{\text{o}})$, forces the learned representation $Z$ to be a minimal sufficient statistic of the input. In our context, this encourages the model to discard non-essential information, which critically includes the noise introduced by the arbitrary locations of missing values. Following the variational inference (Voloshynovskiy et al., 2019), we derive a equivalent form of the compactness term in Eq. 1 as follows:
$I_{\theta}(Z;X^{\text{o}})$ $=\mathbb{E}_{p(x^{\text{o}},z)}[\log\frac{p(x^{\text{o}},z)}{p(z)\cdot p(x^{\text{o}})}]=\mathbb{E}_{p(x^{\text{o}})}[D_{KL}(p(z|x^{\text{o}})||q(z))]-D_{KL}[p(z)\ ||\ q(z)].$ (6)
Because of difficulty in posterior calculation and the non-negative property of Kullback-Leibler (KL) divergence, we use $p_{\theta}(z|x^{\text{o}})$ to approximate the true posterior distribution $p(z|x^{\text{o}})$ and bound Eq. 6:
@@
where $\mu_{\theta}(\cdot)$ and $\sigma_{\theta}(\cdot)$ are designed as neural networks with parameter $\theta$. For training, we use the standard reparameterization trick (Kingma, 2013), $Z=\mu_{\theta}(X^{\text{o}})+\sigma_{\theta}(X^{\text{o}})\odot\epsilon$, which makes the objective in Eq. 7 differentiable without the need for stochastic estimation as follows:
$\mathcal{L}_{\text{Comp}}=\frac{1}{2}\sum_{j=1}^{D}\left(1+\log\left(\sigma_{\theta}^{(j)}(X^{\text{o}})\right)^{2}-\left(\mu_{\theta}^{(j)}(X^{\text{o}})\right)^{2}-\left(\sigma_{\theta}^{(j)}(X^{\text{o}})\right)^{2}\right).$ (9)
@@ 3.4.2 Informativeness Principle
To balance the compactness objective, the informativeness principle ensures that the representation $Z$ preserves sufficient information for the forecasting task. To derive a tractable lower bound for the informativeness term, we follow the framework in (Voloshynovskiy et al., 1912) and Eq. 2, and assume that time series data follow a Gaussian distribution with fixed variance ($\sigma^{2}I$), i.e., $q_{\theta}(y|z)=\mathcal{N}(\widehat{y},\sigma^{2}I)$ (Choi & Lee, 2023). The derivation proceeds as follows:
$I_{\theta}(Y;Z)$ $=\mathbb{E}_{p(z,y)}[\log\frac{p(y|z)}{p(y)}]=\mathbb{E}_{p(z,y)}[\log\frac{q_{\theta}(y|z)}{p(y)}]+\mathbb{E}_{p(z,y)}[\log\frac{p(y|z)}{q_{\theta}(y|z)}],$ (10)
$\geq\mathbb{E}_{p(z,y)}[\log q_{\theta}(y|z)]=\mathbb{E}_{p(z,y)}\left[\frac{1}{2\sigma^{2}}\|y-\widehat{y}\|^{2}+\frac{T}{2}\log(2\pi\sigma^{2})\right],$
$\propto\mathbb{E}_{p(z,y)}\left[\|y-\widehat{y}\|^{2}\right]\stackrel{{\scriptstyle\text{def}}}{{=}}-\mathcal{L}_{\text{Pred}},$

### v3

@@ Abstract
Missing values are common in real-world time series, and multivariate time series forecasting with missing values (MTSF-M) has become a crucial area of research for ensuring reliable predictions. To address the challenge of missing data, current approaches have developed an imputation-then-prediction framework that uses imputation modules to fill in missing values, followed by forecasting on the imputed data. However, this framework overlooks a critical issue: there is no ground truth for the missing values, making the imputation process susceptible to errors that can degrade prediction accuracy. In this paper, we conduct a systematic empirical study and reveal that imputation without direct supervision can corrupt the underlying data distribution and actively degrade prediction accuracy. To address this, we propose a paradigm shift that moves away from imputation and directly predicts based on the partially observed time series. We introduce Consistency-Regularized Information Bottleneck (CRIB), a novel framework built on the Information Bottleneck principle. CRIB combines a unified-variate attention mechanism with a consistency regularization scheme to learn robust representations that filter out noise introduced by missing values while preserving essential predictive signals. Comprehensive experiments on several real-world datasets demonstrate the effectiveness of CRIB, which predicts accurately even under high missing rates. Our code is available in https://github.com/Muyiiiii/CRIB.
Machine Learning, ICML
@@ 1 Introduction
Multivariate time series forecasting (MTSF), which aims to predict future values of multiple variates based on historical observations, plays an important role in many domains, such as traffic flow forecasting (Shang et al., 2022; Yu et al., 2017; Bai et al., 2020), financial analysis (Schaffer et al., 2021; Zhu et al., 2024; Li et al., 2025; Hu et al., 2025b, a), and weather prediction (Zheng et al., 2015; Wu et al., 2021; Tan et al., 2022). However, due to uncontrollable factors such as data collection difficulties and transmission failures (Li et al., 2023; Marisca et al., 2022; Cini et al., 2021; Zhang et al., 2025a), real-world multivariate time series data is often partially observed, with missing values scattered throughout the series. These missing values inevitably introduce noise, leading to distribution shifts and disrupting the variate correlations. MTSF models (Cao et al., 2020; Liu et al., 2022; Ekambaram et al., 2023; Hu et al., 2025e), which typically rely on complete data, are highly sensitive to such shifts and correlation destruction, thus failing to make accurate predictions (Zhou et al., 2023; Hu et al., 2025c). This has driven increasing interest in multivariate time series forecasting with missing values (MTSF-M) (Cao et al., 2018; Zuo et al., 2023; Tang et al., 2020), where the objective is to generate accurate and robust forecasts despite the presence of incomplete data.
@@
To mitigate the impact of missing values, recent MTSF-M research (Yu et al., 2025; Peng et al., 2025) has focused on enhancing observed data by imputing missing values to improve prediction performance. One common approach is the two-stage framework, where an imputation module (Wu et al., 2022; Cao et al., 2018; Du et al., 2023) first fills in the missing values, and a forecasting model then predicts future values based on the imputed data (Peng et al., 2025; Chen et al., 2023; Wu et al., 2015). Moreover, to reduce error accumulation between these two stages of two separate models, some studies have proposed an end-to-end framework (Yu et al., 2024, 2025) that imputes missing values progressively during encoding and performs forecasting using the imputed representations. Overall, these methods generally follow an imputation-then-prediction paradigm, aiming to improve forecasting accuracy by mitigating the negative effects of missing values compared to directly applying forecasting models to incomplete data.
@@
Figure 1: Analysis of the imputation-then-prediction paradigm on PEMS-BAY (40% missing rate). (a) t-SNE visualizations show that current imputation modules cannot recover the original data distribution and their forecasts mismatch with the prediction target, while our direct-prediction method aligns better with the target. (b, c) Correlation maps reveal that imputation fails to recover true variate correlations, whereas our method preserves underlying correlations more effectively. Quantitative analysis is in App. B.
@@
However, current MTSF-M methods ignore a critical limitation in real-world applications: there is no ground truth for missing values. In such scenarios, the imputation module of the current MTSF-M methods would lack reliable guidance, which means the imputed values and reconstructed correlations cannot be guaranteed to be accurate with only the final prediction guidance. As a result, noise would propagate into the prediction stage and degrade forecasting performance, particularly when the missing rate is high. To investigate this issue, we conduct an empirical analysis of representative imputation-then-prediction methods, where original and observed data denote the complete and partially observed data, respectively. This includes the two-stage framework combining TimesNet (Wu et al., 2022) for imputation and TimeXer (Wang et al., 2024b) for forecasting, as well as the end-to-end framework BiTGraph (Chen et al., 2023). Figure 1 illustrates the empirical results, where panel (a) visualizes the distributions of imputed and predicted values, and panels (b) and (c) present the correlations among variates. Our findings highlight two key phenomena:
@@
  • Improper imputation can corrupt the observed data. Current MTSF-M frameworks commonly employ imputation modules to recover missing values. However, as shown in Figure 1 (a-1, b), without enough direct supervision, imputed values deviate considerably from the distribution of the original complete data, and the underlying correlations among variates are not correctly reconstructed. The deterioration of both the data distribution and variate correlations suggests that imputation guided only by the prediction objective can degrade the observed data rather than effectively repairing the missing values.
@@
  • Flawed imputation, in turn, leads to poor prediction performance. Errors from the imputation stage inevitably propagate into forecasting. As shown in Figure 1 (a-2, c), the predictions exhibit large deviations from the prediction targets. Notably, even a model TimeXer applied directly to incomplete observed data outperforms a more complex framework that combines TimesNet for imputation with TimeXer for prediction. These findings indicate that a flawed imputation stage can actively harm, rather than enhance, the forecasting capabilities of a model.
@@
  • Empirical analysis: We perform a systematic empirical study of dominant imputation-then-prediction paradigms for MTSF-M, showing that imputation modules guided only by the prediction objective can corrupt the observed data distribution and degrade prediction performance.
  • Method: We propose a novel direct-prediction method, CRIB, which removes the imputation step completely. Built on the Information Bottleneck principle, CRIB integrates unified-variate attention and consistency regularization to learn refined representations that balance noise filtering and task-relevant information preservation.
  • Experiments: Extensive experiments on four real-world benchmarks demonstrate that CRIB significantly outperforms state-of-the-art methods by an average of 18%, particularly under high missing rates, highlighting the advantage of direct prediction over imputation-then-prediction.
@@ 2 Preliminaries
IB theory (Tishby and Zaslavsky, 2015; Voloshynovskiy et al., 2019; Yang et al., 2025) provides an information-theoretic framework for learning compact yet informative representations. Given the partially observed input $X^{\text{o}}$ and prediction target $Y$, the goal is to learn a latent representation $Z$ that compresses $X^{\text{o}}$ while preserving maximal information about $Y$. This trade-off in CRIB is formalized as follows:
@@
Here, $\theta$ represents the learnable parameters of our proposed CRIB. $I(Z;X^{\text{o}})$ and $I(Y;Z)$ are the mutual information terms measuring compactness and informativeness, respectively. The Lagrange multiplier $\beta\in\mathbb{R}^{+}$ controls the balance between these two terms (Tishby et al., 2000). Furthermore, under standard assumptions in the IB literature (Alemi et al., 2016; Chalk et al., 2016; Ma et al., 2023), the joint distribution of the variables can be factorized as:
$p(X^{\text{o}},Y,Z)$ $=p(Z\mid X^{\text{o}},Y)\,p(Y\mid X^{\text{o}})\,p(X^{\text{o}})$ (2)
$=p(Z\mid X^{\text{o}})\,p(Y\mid X^{\text{o}})\,p(X^{\text{o}}),$
@@ 3 Methodology
Figure 2: Overall framework of CRIB. (a) Data Augmentation creates a more challenging view of the partially observed data $X^{\text{o}}$ by generating an augmented version $X^{\text{Aug}}$. (b) The Patching Embedding layer converts the $X^{\text{o}}$ and $X^{\text{Aug}}$ into robust patch-level feature representations $H$ and $H^{\text{Aug}}$. (c) The Unified-Variate Attention mechanism models the global correlations between all the patches within $H$ and $H^{\text{Aug}}$ to produce refined representations $Z$ and $Z^{\text{Aug}}$. (d) Consistency Regularization aligns the representations from the original $Z$ and the augmented views $Z^{\text{Aug}}$. The entire process is guided by the IB principles of compactness and informativeness to predict $\widehat{Y}$.
As illustrated in Figure 2, our proposed model, CRIB, bypasses the problematic imputation stage by performing forecasts directly on the partially observed data. The architecture is composed of several key stages, each designed to address the challenges of learning from partially observed data. First, to handle the raw, sparse input, we introduce a Patching Embedding layer that employs a Temporal Convolutional Network (TCN) (Bai et al., 2018) to learn robust local feature representations from available data points. Second, to capture the complex global correlations that are disrupted by missingness, a Unified-Variate Attention mechanism models correlations across all patches simultaneously. Third, to ensure the model learns features that are stable and invariant to different missingness, especially under high missing rates, we introduce a Consistency Regularization scheme based on data augmentation. The entire learning process is guided by the IB principle, which provides a theoretical foundation for learning a representation that is maximally compressive against noise while being sufficiently informative for the forecasting task.
@@ 3.1 Patching Embedding
$\text{TE}(t,m)=\begin{cases}\ \sin(t/10000^{2m/P})&\text{ if }m=2k,\\ \ \cos(t/10000^{2m/P})&\text{ if }m=2k+1,\end{cases}$ (3)
@@ 3.4 Information Bottleneck Guidance
To enhance the quality of the learned representations $Z$ and improve forecasting accuracy, we propose an IB-based guidance. This guidance aims to balance compactness (filtering out irrelevant information) with informativeness (preserving relevant task-specific signals), allowing CRIB to focus on the most significant factors for accurate forecasting. In this section, we present how the compactness and informativeness principles are formulated and implemented in our framework. Full derivations are detailed in App. A.
@@ 3.4.1 Compactness Principle
The compactness principle, which aims to minimize the mutual information $I_{\theta}(Z;X^{\text{o}})$, forces the learned representation $Z$ to be a minimal sufficient statistic of the input. In our context, this encourages the model to discard non-essential information, which critically includes the noise introduced by the arbitrary locations of missing values. Following the variational inference (Voloshynovskiy et al., 2019), we derive a equivalent form of the compactness term in Equation 1:
$I_{\theta}(Z;X^{\text{o}})$ $=\mathbb{E}_{p(x^{\text{o}},z)}\!\left[\log\frac{p(x^{\text{o}},z)}{p(z)\,p(x^{\text{o}})}\right]$ (6)
$=\mathbb{E}_{p(x^{\text{o}})}\!\left[D_{\mathrm{KL}}\!\left(p(z\mid x^{\text{o}})\,\|\,q(z)\right)\right]$
$\quad-D_{\mathrm{KL}}\!\left(p(z)\,\|\,q(z)\right).$
Because of difficulty in posterior calculation and the non-negative property of Kullback-Leibler (KL) divergence, we use $p_{\theta}(z|x^{\text{o}})$ to approximate the true posterior distribution $p(z|x^{\text{o}})$ and bound Equation 6:
@@
where $\mu_{\theta}(\cdot)$ and $\sigma_{\theta}(\cdot)$ are designed as neural networks with parameter $\theta$. For training, we use the standard reparameterization trick (Kingma, 2013), $Z=\mu_{\theta}(X^{\text{o}})+\sigma_{\theta}(X^{\text{o}})\odot\epsilon$, which makes the objective in Equation 7 differentiable without the need for stochastic estimation as follows:
$\mathcal{L}_{\text{Comp}}$ $=\frac{1}{2}\sum_{j=1}^{D}\Big(1+\log\big(\sigma_{\theta}^{(j)}(X^{\text{o}})\big)^{2}$ (9)
$\qquad-\big(\mu_{\theta}^{(j)}(X^{\text{o}})\big)^{2}-\big(\sigma_{\theta}^{(j)}(X^{\text{o}})\big)^{2}\Big).$
@@ 3.4.2 Informativeness Principle
To balance the compactness objective, the informativeness principle ensures that the representation $Z$ preserves sufficient information for the forecasting task. To derive a tractable lower bound for the informativeness term, we follow the framework in (Voloshynovskiy et al., 1912) and Equation 2, and assume that prediction errors follow a Gaussian distribution with fixed variance $\sigma^{2}$, i.e., $q_{\theta}(y|z)=\mathcal{N}(\widehat{y},\sigma^{2}I)$ (Choi and Lee, 2023). The derivation proceeds as:
$I_{\theta}(Y;Z)$ $=\mathbb{E}_{p(z,y)}\left[\log\frac{q_{\theta}(y|z)}{p(y)}\right]+\mathbb{E}_{p(z,y)}\left[\log\frac{p(y|z)}{q_{\theta}(y|z)}\right]$ (10)
$\geq\mathbb{E}_{p(z,y)}[\log q_{\theta}(y|z)]$
$=-\mathbb{E}_{p(z,y)}\left[\frac{1}{2\sigma^{2}}\|y-\widehat{y}\|^{2}+\frac{T}{2}\log(2\pi\sigma^{2})\right]$
$\propto-\mathbb{E}_{p(z,y)}\left[\|y-\widehat{y}\|^{2}\right]\stackrel{{\scriptstyle\text{def}}}{{=}}-\mathcal{L}_{\text{Pred}}.$